# JWST Reveals Varied Origins Between Jupiter’s Irregular Satellites

### v1

@@ Abstract
We report observations of eight Jovian irregular satellites with JWST’s NIRSpec instrument: Himalia, Elara, Pasiphae, Sinope, Lysithea, Carme, Ananke, and Themisto. Irregular satellite families, which are presumed to have formed via collisions, contain various Trojan-like and C-type-asteroid-like surfaces. We sample the three largest members of the Himalia satellite family, detecting the presence of complexed CO2 and a unique absorption band from whose character correlates with satellite size. The two largest irregular satellites, Himalia family members Himalia and Elara, contain ammoniated phyllosilicates that are not seen in the meteorite inventory. We propose that the Himalia parent body was heterogeneous and formed with materials similar to Ceres-like ammonium-bearing asteroids. Several small () irregular satellites closely track the colors and absorption bands of “red” Jovian Trojans, demonstrating that these compositions are retained amongst the products of collisions that occurred after Jovian capture. We report the first detection of aqueous alteration products in the retrograde satellite swarm, finding Ananke’s 3 band to closely match phyllosilicates seen in C2 chondrites. Notably, objects with OH absorption features similar to the Trojan asteroid Eurybates are found in both the retrograde Pasiphae family and the prograde Himalia family, confounding a simple link between such materials and a single surface type. The irregular satellites appear consistent with some materials that experienced alteration from liquid water and others that did not. Consequently, Jupiter may have captured bodies that formed from different initial compositions, or bodies that experienced different levels of heating, driving differential alteration processes.
@@ 1 Jupiter’s Irregular Satellites: Possible Remnants of Outer Solar System Mixing
The origins of Jupiter’s irregular satellites are uncertain. The highly inclined and eccentric orbits of these small satellites lie exterior to the regular, Galilean satellites, suggesting that they were captured from asteroid- or comet-like parent reservoirs (e.g., Sheppard & Jewitt, 2003; Nesvorný et al., 2007, 2014). Jovian irregular satellite orbits are also clustered, with one large prograde group (Himalia), three major retrograde groups (Ananke, Carme, and Pasiphae), and several smaller, isolated satellites (Sheppard & Jewitt, 2003). The (a,i) orbital elements of the Jovian irregular satellites, noting major family clusters, are displayed in Figure 1.
Various processes have been proposed to explain the capture of Jupiter’s irregular satellites, as well as similar populations at Saturn, Uranus, and Neptune. Nesvorný (2018), Jewitt & Haghighipour (2007), and Nicholson et al. (2008) provide thorough reviews of these mechanisms. Some relate the capture process to the formation of the planet itself, including pull-down capture (Heppenheimer & Porco, 1977, where the planet’s mass growth captures nearby resonant objects) and gas-drag (Pollack et al., 1979; Astakhov et al., 2003; Ćuk & Burns, 2004; Kortenkamp, 2005, where nebular gas affects nearby planetesimals that drift into bound states). Alternatively, three-body interactions could emplace satellites after planet formation (Colombo & Franklin, 1971).
@@
Through the lens of migratory capture, the most likely source reservoir for irregular satellites and Trojans would be the primordial Kuiper belt. This is a hypothesized region stretching from and au (Nesvorný et al., 2019), immediately outside the locations where the giant planets accreted ( and au, e.g., Tsiganis et al., 2005; Nesvorný & Morbidelli, 2012; Nesvorný et al., 2019). Within Myr after the dissipation of the solar nebula, Neptune’s migration through the primordial Kuiper belt would have destabilized this planetesimal population and triggered dynamic instability throughout the outer solar system (Tsiganis et al., 2005). Planetesimals ejected inwards would enter planet-crossing orbits, and those ejected outwards would create the scattered disk (Levison et al., 2008). Meanwhile, the present-day dynamically ”cold” Kuiper belt population would represent the portion of the primordial Kuiper belt that formed far enough away from the giant planets to remain stable (although this population may have also been largely depleted in mass, see e.g. Gomes et al., 2018)
@@
While the irregular satellites could have been captured from a single source, they could also represent a composite of various populations that interacted with Jupiter over the planet’s history. The three-body capture scenarios defined by Nesvorný et al. (2007) occur in the vicinity of primordial Kuiper belt remnants, not the original generation of middle solar system planetesimals. However, some closely-bound irregular satellites may be able to survive from previous capture events. For example, Nesvorný et al. (2014) find that of Himalia-like Jovian satellites could survive planetary encounters that occurred after their capture. The likelihood of such a scenario is currently speculative, but can be constrained by determining the variety of the present-day irregular satellite population. Determining whether or not all irregular satellite parent bodies formed within similar primordial environments provides a key constraint on the character of Jupiter’s migratory history.
@@
Observational evidence does not yet uniquely discriminate irregular satellites’ formation region(s). Jupiter’s position, proximate to both the inner solar system’s reservoirs of asteroids as well as the outer solar system’s comets, trans-Neptunian objects (TNOs), and Centaurs, provides a rich array of comparison populations. Determining the irregular satellites’ compositions in the context of these other populations tests how migratory capture could have occurred. But the post-capture evolution of the irregular satellites, including surface alteration and collisional breakup, must be accounted for to forge successful cross-population links. To determine the lineage(s) of Jupiter’s captured satellites, we therefore require knowledge of their bulk compositions and the variability within various collisional families. We provide these bulk compositional assessments with high-SNR JWST spectra of the largest objects in each each major satellite family, and we provide the first detailed exploration of variability within a family as a function of size. These constraints address irregular satellite formation conditions and the evolutionary processes that have shaped them.
@@ 1.1 Previous Observational and Laboratory Context
Numerous spacecraft missions have explored the Jovian system, but only the largest irregular satellite, Himalia, has been the subject of resolved spacecraft imaging (Porco et al., 2003, resolved across 5 pixels) or spectroscopy (Brown et al., 2003, SNR). Telescopic observations have been the primary method to understand the population. These characterization efforts suggest kinship between irregular satellites and various groups of dark primitive asteroids. Brown & Rhoden (2014) found that Himalia has a spectral feature that indicates the presence of hydrated material and appears similar to the main belt asteroid (52) Europa (Takir & Emery, 2012). This feature is only partially observable from ground-based telescopes, which limits compositional tests beyond that comparison. Himalia and its associated family were also reported to display a 0.7 feature of variable strength correlated with size (Vilas & Hendrix, 2024).
Amongst other satellite families, numerous neutral-to-very-red spectra have been observed, similar in color to Jupiter Trojan asteroids and without apparent absorption features at wavelengths shorter than (Sharkey et al., 2023; Bhatt et al., 2017). Coarser color photometry surveys (Graykowski & Jewitt, 2018; Grav et al., 2003; Grav & Holman, 2004) have characterized a wider satellite sample and have found Jupiter’s irregular satellites to have colors similar to those of Saturn, Uranus, and Neptune. But the relative scarcity and faintness of Jupiter’s irregular satellites have limited further attempts to characterize them. These observational constraints generally limit even photometric studies to the few Jovian satellites that are larger than D 10 km, and they have precluded the high-SNR spectroscopy that is needed to interpret mineral absorption features.
Existing high-resolution, high-SNR spectra of Galilean satellites may provide an alternate means to investigate irregular satellite compositions. Collisions within irregular satellite swarms generates dust that likely collides with the surfaces of the Galilean moons, in particular the leading hemisphere of Callisto (e.g., Bottke et al., 2010, 2013, 2024; Chen et al., 2024). Accumulated dust from these primordial bodies then mixes with native components in the Galilean moons’ regoliths that is subsequently irradiated by charged particles trapped in Jupiter’s massive magnetosphere (Johnson et al., 2004). This process would form new components, such as CO2 (e.g., Gomis & Strazzulla, 2005), sulfates (e.g., Carlson et al., 2002), and perhaps complex CN-bearing organics (e.g., Cartwright et al., 2024) depending on the nature of the exogenic material.
@@
With the advent of JWST, our understanding of small solar system bodies is rapidly expanding. JWST’s NIRSpec instrument’s high sensitivity over key infrared passbands identifies assemblages of rocky minerals, volatile ices, and organic compounds with spectroscopic complexity previously accessible only via laboratory study. Recent observations of Jovian Trojans by Wong et al. (2024) highlight the importance of studying carbonaceous asteroids from 0.7-5.2 . When observed from Earth-based telescopes, Trojans are featureless 2.5 . When observed via JWST, distinct spectral features from 2.5-5.0 in both color groups were discovered, providing the first identifications of organic compounds and OH-bearing materials on these asteroids.
@@
Coupled with their shared heliocentric distance, the strong spectral slope similarities between the Jovian Trojans and irregular satellites already suggest some level of compositional similarity. However, the effects of more frequent collisions within the irregular satellites (Bottke et al., 2010), and the presence of objects similar to outer-main belt asteroids (Brown & Rhoden, 2014; Takir & Emery, 2012) complicates a direct equivalence between the two populations. This work details the mineralogies of Jupiter’s irregular satellite groups, shedding new light on the organic-rich material reservoirs that have interacted with Jupiter and the Galilean moons.
@@ 2 Observations and Data Reduction
Satellite | Diameter (km) | Observation Date | Phase Angle | Sun-Target | Target-Observer |
(Degrees) | Distance (au) | Distance (au) | Distance (au) | ||
JXVIII Themisto | 8aaVia Sheppard & Jewitt (2003), assumes geometric albedo of 0.04. | 2024-01-31 | 11.59 | 4.94 | 4.90 |
JVI Himalia | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-31 | 11.54 | 4.96 | 4.92 |
JX Lysithea | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-25 | 11.55 | 4.96 | 4.82 |
JVII Elara | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-18 | 11.40 | 4.97 | 4.72 |
JXII Ananke | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-31 | 11.19 | 5.11 | 5.08 |
JXI Carme | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-25 | 11.92 | 4.81 | 4.69 |
JVIII Pasiphae | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-25 | 11.53 | 4.97 | 4.89 |
JIX Sinope | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-21 | 11.58 | 4.94 | 4.76 |
Observations were performed using the NIRSpec integral field unit (IFU) in prism mode () for all targets except Himalia, which was observed in the higher resolution G235M and G395M grating modes (). Observational circumstances, including timing and geometry, are given in Table 1.
@@
Spectral extraction on each individual data cube made use of the “template PSF-fitting” routine first described in Wong et al. (2024). The uniformly spaced wavelength array was constructed using the CRVAL3 and CDELT3 header keywords and slices 0.7 m and 5.1 were trimmed due to higher noise. Then, the DQ extension was used to identify pixels flagged by the pipeline and any pixel with a non-zero value was set to NaN for the remainder of the extraction. Target centroids were first determined by-eye (typically good to a few tenths of a pixel) and used to measure radial distances from the target. Outlier pixels due to artifacts or background objects, as well as those 5 pixels from the centroid were set to NaN. Then the background was calculated in each slice of the cube using the median of all pixels 5 pixels from the centroid and subtracted from that slice. A template PSF was then computed as the median of the 10 background-subtracted slices proceeding and succeeding the slice under consideration (inclusive) for a total of 21 slices. Slices near the edge of the considered wavelength range included 21 total slices, but the resulting template PSFs were still adequate. The template PSF was then trimmed to a 99-pixel vignette, with all external pixels set to NaN, in order to reduce the leverage of more distant pixels when fitting to the central slice in the set. The integrated flux of the vignette was then normalized to unity. The scipy.optimize.minimize function and a Nelder-Mead (amoeba) algorithm was used to determine the best-fit values for the flux scaling factor of the template PSF and the background level. This fit was iteratively performed 3 times, using the outputs from the previous iteration as the initial guess for the next iteration. Additionally, after each iteration a model was constructed (template PSF multiplied by the flux scaling factor followed by the addition of the background) and subtracted from the data, with any pixels 5- from the mean counted as outliers and set to NaN prior to the next iteration. After the third iteration, a center-of-mass centroid was computed to determine the target centroid to sub-pixel accuracy, the template PSF was multiplied by the flux scaling factor, and the flux was extracted within a 3.5-pixel radius circular aperture.

### v2

@@ Abstract
We report JWST NIRSpec (0.7-5.1 ) observations of eight Jovian irregular satellites across five orbital groups. We detect variation in the phyllosilicate content of the the three largest members of the Himalia collisional family (Himalia, , Elara, , and Lysithea, ). Himalia contains complexed CO2 and overlapping absorption features from that match laboratory samples of ammoniated phyllosilicates. Lysithea displays a simpler, single-minimum feature caused by an unidentified absorber. Elara presents a 3 band that matches by a simple average of Himalia and Lysithea. We argue that the Himalia parent body was heterogeneous and formed with materials similar to Ceres-like ammonium-bearing asteroids, consistent with suggestions from previous visible-wavelength observations. The satellites Carme, Sinope, and Themisto have colors and absorption features similar to “red” Jovian Trojans. The satellites Ananke, Pasiphae, and Lysithea each have absorption bands centered from , intermediate between the longer-wavelength bands observed on Trojan asteroids and the shorter-wavelength bands commonly seen on phyllosilicate-rich C2 chondrites. These intermediate-wavelength absorptions are present in both Trojan-like families and the hydrated Himalia family, confounding a link to a single compositional context. The observed differences between Jupiter’s irregular satellites and Trojans requires one of two possibilities: 1) some Trojan parent bodies contained hydrated materials in their cores; 2) some Jovian irregular satellites were not captured from the Trojan parent reservoir. Future searches for NH-bearing materials across the solar system and 3 studies of small body collisional families may provide means to discriminate between these two hypotheses.
@@ 1 Jupiter’s Irregular Satellites: Possible Remnants of Outer Solar System Mixing
The origins of Jupiter’s irregular satellites are uncertain. The highly inclined and eccentric orbits of these small satellites lie exterior to the regular, Galilean satellites, suggesting that they were captured from asteroid- or comet-like parent reservoirs (e.g., Sheppard & Jewitt, 2003; Nesvorný et al., 2007, 2014). Jovian irregular satellite orbits are also clustered, with one large prograde group (Himalia), three major retrograde groups (Ananke, Carme, and Pasiphae), and several smaller, isolated satellites (Sheppard & Jewitt, 2003). The (a,i) orbital elements of the 95 known Jovian irregular satellites (Sheppard et al., 2023), noting major family clusters, are displayed in Figure 1.
Various processes have been proposed to explain the capture of Jupiter’s irregular satellites, as well as similar populations at Saturn, Uranus, and Neptune. Nesvorný (2018), Jewitt & Haghighipour (2007), and Nicholson et al. (2008) provide thorough reviews of these mechanisms. Some relate the capture process to the formation of the planet itself, including pull-down capture (Heppenheimer & Porco, 1977, where the planet’s mass growth captures nearby resonant objects) and gas-drag (Pollack et al., 1979; Astakhov et al., 2003; Ćuk & Burns, 2004; Kortenkamp, 2005, where nebular gas affects nearby planetesimals that drift into bound states). Alternatively, three-body interactions could emplace satellites after planet formation (Colombo & Franklin, 1971), including recent suggestions that close stellar flybys could have implanted the irregular satellites (Pfalzner et al., 2024).
@@
Through the lens of migratory capture, the most likely source reservoir for irregular satellites and Trojans would be the primordial Kuiper belt. This is a hypothesized region stretching from and au (Nesvorný et al., 2019), immediately outside the locations where the giant planets accreted ( and au, e.g., Tsiganis et al., 2005; Nesvorný & Morbidelli, 2012; Nesvorný et al., 2019). Within Myr after the dissipation of the solar nebula, Neptune’s migration through the primordial Kuiper belt would have destabilized this planetesimal population and triggered dynamic instability throughout the outer solar system (Tsiganis et al., 2005). Planetesimals ejected inwards would enter planet-crossing orbits, and those ejected outwards would create the scattered disk (Levison et al., 2008). Meanwhile, the present-day dynamically “cold” Kuiper belt population would represent the portion of the primordial Kuiper belt that formed far enough away from the giant planets to remain stable (although this population may have also been largely depleted in mass, see e.g. Gomes et al., 2018)
@@
While the irregular satellites could have been captured from a single source, they could also represent a composite of various populations that interacted with Jupiter over the planet’s history. The three-body capture scenarios defined by Nesvorný et al. (2007) occur in the vicinity of primordial Kuiper belt remnants, not the original generation of middle solar system planetesimals. However, some closely-bound irregular satellites may be able to survive from previous capture events. For example, Nesvorný et al. (2014) find that of Jovian satellites with closer-in, Himalia-like orbits could survive planetary encounters that occurred after their capture. The likelihood of this successive capture scenario is currently speculative, but could be constrained by determining whether or not all irregular satellite parent bodies formed within similar primordial environments.
@@
Observational evidence does not yet uniquely discriminate irregular satellites’ formation region(s). Jupiter’s position, proximate to both the inner solar system’s reservoirs of asteroids as well as the outer solar system’s comets, trans-Neptunian objects (TNOs), and Centaurs, provides a rich array of comparison populations. Determining the irregular satellites’ compositions in the context of these other populations tests how migratory capture could have occurred. But the post-capture evolution of the irregular satellites, including surface alteration and collisional breakup, must be accounted for in order to forge successful cross-population links. To determine the lineage(s) of Jupiter’s captured satellites, we therefore require knowledge of their bulk compositions as well as the variability within various collisional families. We provide these bulk compositional assessments with high-SNR JWST spectra of the largest objects in each major satellite family, and we provide the first detailed exploration of variability within a collisional family as a function of size. These constraints address irregular satellite formation conditions and the evolutionary processes that have shaped them.
@@ 1.1 Previous Observational and Laboratory Context
Numerous spacecraft missions have explored the Jovian system, but only the largest irregular satellite, Himalia, has been characterized by spacecraft imaging (resolved across 5 pixels by Porco et al. 2003, unresolved by Cheng et al. 2010) and spectroscopy (Brown et al., 2003; Chamberlain & Brown, 2004, SNR). Chamberlain & Brown (2004) reported the “suggestion of an absorption” near 3 on Himalia, but additional spectral structure could not be retrieved from those data. Telescopic observations have suggested kinship between irregular satellites and various groups of dark primitive asteroids. Jarvis et al. (2000) reported a 0.7 feature on Himalia, suggesting that Himalia contains phyllosilicates and is taxonomically similar to the main belt asteroid (52) Europa. Brown & Rhoden (2014) detected and characterized Himalia’s 3 band, reporting this feature to also match (52) Europa (Takir & Emery, 2012) but not match any known phyllosilicates. This feature is only partially observable from ground-based telescopes because of atmospheric opacity in the region, which limits compositional tests by obscuring diagnostic information like the locations of band centers. Himalia and its associated family were also reported to display a 0.7 feature of variable strength correlated with size by Vilas & Hendrix (2024), who also report a similar feature in some Ananke and Pasiphae spectra. Phyllosilicates are also components of the best spectral fits to Himalia and Elara by Bhatt et al. (2017).
Amongst other satellite families, numerous neutral-to-very-red spectra have been observed, similar in color to Jupiter Trojan asteroids and without apparent absorption features at wavelengths shorter than (Sharkey et al., 2023; Bhatt et al., 2017). Coarser color photometry surveys (Graykowski & Jewitt, 2018; Grav et al., 2003; Grav & Holman, 2004) have characterized a wider satellite sample and have found Jupiter’s irregular satellites to have colors similar to those of Saturn, Uranus, and Neptune. But the relative scarcity and faintness of Jupiter’s irregular satellites have limited further attempts to characterize them. These observational constraints generally limit even photometric studies to the few Jovian satellites that are larger than D 10 km, and they have precluded the high-SNR spectroscopy that is needed to interpret near-infrared mineral absorption features.
Existing high-resolution, high-SNR spectra of Galilean satellites may provide an alternate means to investigate irregular satellite compositions. Collisions within irregular satellite swarms generate dust that likely collides with the surfaces of the Galilean moons, in particular the leading hemisphere of Callisto (e.g., Bottke et al., 2010, 2013, 2024; Chen et al., 2024). Accumulated dust from these primordial bodies then mixes with native components in the Galilean moons’ regoliths that is subsequently irradiated by charged particles trapped in Jupiter’s massive magnetosphere (Johnson et al., 2004). This process would form new components, such as CO2 (e.g., Gomis & Strazzulla, 2005), sulfates (e.g., Carlson et al., 2002), and perhaps complex CN-bearing organics (e.g., Cartwright et al., 2024) depending on the nature of the exogenic material.
@@
With the advent of JWST, our understanding of small solar system bodies is rapidly expanding. JWST’s NIRSpec instrument’s high sensitivity over key infrared passbands identifies assemblages of rocky minerals, volatile ices, and organic compounds with spectroscopic complexity previously accessible only via laboratory study.
Recent observations of Jovian Trojans by Wong et al. (2024) highlight the importance of studying carbonaceous asteroids from 0.7-5.2 . When observed from Earth-based telescopes, Trojans are featureless 2.5 , and the presence of the 3 bands were partially detectable (Brown, 2016) but not fully characterized due to telluric contamination. When observed via JWST, distinct spectral features from 2.5-5.0 in both color groups were characterized in full, providing clear identifications of organic compounds and OH-bearing materials on these asteroids.
@@
Coupled with their shared heliocentric distance, the strong spectral slope similarities between the Jovian Trojans and irregular satellites already suggest some level of compositional similarity. However, the effects of more frequent collisions within the irregular satellites (Bottke et al., 2010), and the presence of objects like Himalia with absorption features that are comparable to outer-main belt asteroids (Jarvis et al., 2000; Brown & Rhoden, 2014; Takir & Emery, 2012; Vilas & Hendrix, 2024) complicates a direct equivalence between the two populations. This work further constrains the mineralogies of Jupiter’s irregular satellite groups, shedding new light on the organic-rich material reservoirs that have interacted with Jupiter and the Galilean moons.
@@ 2 Observations and Data Reduction
| Satellite | Family | Diameter (km) | Observation Date | Phase Angle | Sun-Target | Target-Observer |
| (Degrees) | Distance (au) | Distance (au) | ||||
| JXVIII Themisto | Themisto | 8aaVia Sheppard & Jewitt (2003), assumes geometric albedo of 0.04. | 2024-01-31 | 11.59 | 4.94 | 4.90 |
| JVI Himalia | Himalia | bbVia (Grav et al., 2015) from NEOWISE thermal-IR observations. | 2024-01-31 | 11.54 | 4.96 | 4.92 |
| JX Lysithea | Himalia | bbVia (Grav et al., 2015) from NEOWISE thermal-IR observations. | 2024-01-25 | 11.55 | 4.96 | 4.82 |
| JVII Elara | Himalia | bbVia (Grav et al., 2015) from NEOWISE thermal-IR observations. | 2024-01-18 | 11.40 | 4.97 | 4.72 |
| JXII Ananke | Ananke | bbVia (Grav et al., 2015) from NEOWISE thermal-IR observations. | 2024-01-31 | 11.19 | 5.11 | 5.08 |
| JXI Carme | Carme | bbVia (Grav et al., 2015) from NEOWISE thermal-IR observations. | 2024-01-25 | 11.92 | 4.81 | 4.69 |
| JVIII Pasiphae | Pasiphae | bbVia (Grav et al., 2015) from NEOWISE thermal-IR observations. | 2024-01-25 | 11.53 | 4.97 | 4.89 |
| JIX Sinope | Pasiphae | bbVia (Grav et al., 2015) from NEOWISE thermal-IR observations. | 2024-01-21 | 11.58 | 4.94 | 4.76 |
Observations were performed using the NIRSpec integral field unit (IFU) in prism mode () for all targets except Himalia, which was observed in the higher resolution G235M and G395M grating modes ()111Grating-specific dispersion and spectral resolution files are available at NIRSpec Dispersers and Filters in the JWST documentation.. Observational circumstances, including timing and geometry, are given in Table 1.
@@
Spectral extraction on each individual data cube made use of the “template PSF-fitting” routine first described in Wong et al. (2024). The uniformly spaced wavelength array was constructed and slices 0.7 m and 5.1 m were trimmed due to higher noise. Then, the data quality (DQ) extension was used to identify pixels flagged by the pipeline and any pixel with a non-zero value was set to NaN for the remainder of the extraction. Target centroids were first determined by-eye (typically good to a few tenths of a pixel) and used to measure radial distances from the target. Outlier pixels due to artifacts or background objects, as well as those 5 pixels from the centroid were set to NaN. Then the background was calculated in each slice of the cube using the median of all pixels 5 pixels from the centroid and subtracted from that slice. A template PSF was then computed as the median of the 10 background-subtracted slices proceeding and succeeding the slice under consideration (inclusive) for a total of 21 slices. Slices near the edge of the considered wavelength range included 21 total slices, but the resulting template PSFs were still adequate. The template PSF was then trimmed to a 99-pixel vignette, with all external pixels set to NaN, in order to reduce the leverage of more distant pixels when fitting to the central slice in the set. The integrated flux of the vignette was then normalized to unity. The scipy.optimize.minimize function and a Nelder-Mead (amoeba) algorithm was used to determine the best-fit values for the flux scaling factor of the template PSF and the background level. This fit was iteratively performed 3 times, using the outputs from the previous iteration as the initial guess for the next iteration. Additionally, after each iteration a model was constructed (template PSF multiplied by the flux scaling factor followed by the addition of the background) and subtracted from the data, with any pixels 5- from the mean counted as outliers and set to NaN prior to the next iteration. After the third iteration, a center-of-mass centroid was computed to determine the target centroid to sub-pixel accuracy, the template PSF was multiplied by the flux scaling factor, and the flux was extracted within a 3.5-pixel radius circular aperture.